2 g of impure CaCO3 reacts with HCI to produce 410 mL of CO2 at 1 atmospheric pressure and 27°C. Calculate the percentage purity of the CaCO3 used.

Grade:12th pass

1 Answers

Atmik Goswami
3 years ago
Firstly we convert volume of COat STP.
1×410/300 = 1×V/273
or, V = 373.1 mL
We know at STP
22400 ml of CO2 weighs 44 g
So, 373.1 mL of COweighs (44×373.1)/22400
or weight of COis 0.732875 g.
Now we apply POAC on 'C':
1×x/100= 1×0.732875/44
or x=1.665625 g
So, percentage purity of CaCO3 is = 1.665625/2 ×100=83.28 %
or x=1.665625 g
So percentage of